OpenStudy (anonymous):

(x/7)+(y/8)= 10 solve for y; show steps if possible.(:

OpenStudy (anonymous):

ok. u want y alone. start by moving the x/7 away from it. you should get it from there

OpenStudy (anonymous):

so y/8=10-x/7?

OpenStudy (anonymous):

then multiply by 8?

OpenStudy (anonymous):

....sigh. 8x + 7y = 8*7*10

OpenStudy (anonymous):

this is standard form. then just do y = [-8x - 560]/7 this is now more or less slope form

OpenStudy (anonymous):

yes brit. no ronnc

OpenStudy (anonymous):

80-8x/7=y

OpenStudy (anonymous):

eh.. no way. x/7 + y/8 = 10 so 8x/7 + y = 80 8x + 7y = 560 7y = -8x + 560 OR 560 -8x (same thing) y = [-8x+560]/7 OR [560-8x]/7 (same thing again).. you sleepy tutor?

OpenStudy (anonymous):

you are correct ronnc, but you went the long way to scratch your butt.

OpenStudy (anonymous):

true. but its depends how you want to do it. it can be arguable that I show steps better.

OpenStudy (anonymous):

but you didnt simplify.

OpenStudy (anonymous):

x+5=6/x x(x+5)=6 x^2+5x=6 x^2+5x-6=0 now it is in quadratic equation form x^2-x+6x-6=0 now takes the common x(x-1)+6(x-1)=0 again (x-1) is common (x-1)(x+6)=0 now the multiplication of two terms willbe 0 only when any one of them is 0 so if x-1=0 then x=1 and if x+6=0 then x=-6 so ans is x=1,-6 are you satisfied ??